[figure: 9.7 cm]

MAGNETIC PROPERTIES AND GMR OF SPUTTERED Permalloy/Au MULTILAYERS

B. Szymański and F. Stobiecki Institute of Molecular Physics, Polish Academy of Sciences, Poznań, Poland

Permalloy/Au multilayers are among the most promising materials from the application point of view as they show the highest GMR field sensitivity. However, magnetic and GMR properties of Py/Au multilayers strongly depend on the deposition techniques. In this work we present properties of a series of Mls prepared with the double face to face sputtering technique. The Py = Ni_{83}Fe_{17} sublayer thickness has been kept constant within each film and amounts to 2.4 nm. Gold sublayer has been deposited in a narrow wedge form and its mean thickness varied from 1 to 3.5 nm. The existence of a well defined periodic structure was confirmed by X-ray diffraction. The samples were studied by vibrating sample magnetometer, longitudinal Kerr effect and GMR effect measurements. The results can be summarized as follows: i) the Mls are magnetically soft with H_{c} ≤ 1 Oe and show uniaxial anisotropy with H_{k} ≈ 10 Oe, ii) for d_{Au} close to 1.1 nm the antiferromagnetic exchange coupling is present in a very narrow Au thickness range (∼0.2 nm), iii) the relatively small GMR value (1.5% at RT) is caused by nonperfect AF coupling. Temperature dependences of the magnetization reversal and the GMR effect are discussed.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : B. Szymański

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences PL 60-179 Poznań ul. Smoluchowskiego 17, Poland

Email address : szyman@ifmpan.poznan.pl